# Adaptive Convolution for CNN-based Speech Enhancement Models

### v1

@@ VI-A2 Number of candidate kernels
Adaptive 22449.07 5564.01 15.208 0.777 2.526 3.027 3.309 4.053
@@ VI-A3 Generalization capability of adaptive convolution
However, as the model size increases, the improvements introduced by adaptive convolution become less pronounced. For DPCRN-middle, it achieves approximately a 0.06 PESQ gain and a 0.03 DNSMOS-OVRL gain. While still significant, these improvements are relatively smaller compared to those observed in lightweight models. For DPCRN-large, the improvements are minimal, with only about a 0.2 dB SI-SNR gain, a 0.02 PESQ gain, and a 0.02 DNSMOS-OVRL gain. The diminishing improvement can be attributed to the extensive parameter space inherent in larger models. Specifically, the convolutional kernels in such models already include filters capable of adapting to diverse spectral features, reducing the benefit of additional candidate kernels. Moreover, the exponential increase in parameters introduced by adaptive convolution may make training more challenging, further limiting its effectiveness in larger models.
For larger DCCRN model, however, the improvements introduced by adaptive convolution are more pronounced and comparable to those observed in DPCRN-middle. This can be attributed to DCCRN’s relatively lower baseline performance, which suggests a smaller effective parameter space. As a result, adaptive convolution yields more significant gains in this context. Moreover, since convolution operations account for over 97% of DCCRN’s MACs, the proposed strategy, which focuses on optimizing CNNs, becomes even more effective.
@@ VI-B Experimental Results for AdaptCRN
TABLE IV: Ablation study results for AdaptCRN on DNS5 test set.
  Model Para. (K) MACs (M) SI-SNR ESTOI PESQ DNSMOS
OVRL SIG BAK
Noisy - - 5.366 0.625 1.391 2.147 2.918 2.340
AdaptCRN 134.51 37.92 14.892 0.759 2.387 2.939 3.219 4.037
w/o Joint channel attention 124.68 37.30 14.718 0.756 2.352 2.925 3.207 4.031
w/o Joint multi-layer kernel attention 169.21 40.31 14.803 0.759 2.386 2.939 3.221 4.036
w/o Dynamic range compression 134.51 37.92 14.695 0.755 2.342 2.929 3.211 4.031
DW(8)-PW(8)-skip-PW(8) 134.51 37.13 14.672 0.755 2.358 2.931 3.212 4.034
DW(8)-PW(8) 112.88 30.32 14.551 0.754 2.343 2.926 3.207 4.031
DW(8)-PW(64) 255.00 31.52 14.826 0.759 2.390 0.942 3.221 4.039
@@
AdaptCRN 134.51 37.92 14.892 0.759 2.387 2.939 3.219 4.037
@@ VI-B1 Ablation study results for AdaptCRN
We evaluate the effectiveness of the proposed structures in AdaptCRN through ablation experiments by selectively removing specific modules. The results, summarized in Table IV, highlight the pivotal role of adaptive convolution in enhancing model performance. With an increase in computational complexity of less than 5M MACs, adaptive convolution delivers significant improvements, including nearly 1 dB in SI-SNR, 0.02 in ESTOI, 0.19 in PESQ, and over 0.06 in DNSMOS-OVRL. In the proposed model, we utilize adaptive convolution with $K=8$, as this configuration provides a favorable balance between performance gains and parameter amount. Due to the extremely low parameter count of the basic model, the additional parameters introduced by adaptive convolution remain within an acceptable range for real-time applications.
@@
Table IV further presents the model’s performance when the activation function between the two PW Conv layers in the adaptive block is removed (DW(8)-PW(8)-skip-PW(8)), where the numbers in parentheses (e.g., 8) denote $K$, the number of candidate kernels. This modification results in some performance degradation but still outperforms a single PW Conv layer with $K=8$ (DW(8)-PW(8)). As analyzed in Section IV-C, the former configuration can be reparameterized into a single layer with 64 candidate kernels, showcasing the enhanced nonlinear representational capacity of adaptive convolution. Moreover, directly optimizing 64 independent kernels (DW(8)-PW(64)) achieves better performance compared to the reparameterized result, due to the lower-rank kernel space of the latter. This configuration even surpasses the original AdaptCRN, demonstrating adaptive convolution’s potential to significantly improve performance while lowering computational overhead, albeit with an increase in parameter amount, as discussed in Section IV-C and Section VI-A2.
@@ VI-B2 Comparison with the baseline models
We evaluate our ultra-lightweight model, AdaptCRN, against DPCRN-light, GTCRN, and LiSenNet on DNS5 test set. As illustrated in Table V, AdaptCRN outperforms the baseline models across all metrics. AdaptCRN remains efficient with computational complexity fewer than 40M MACs, which is only slightly higher than GTCRN. While it does have more parameters compared to the baselines, it remains within acceptable limits for practical applications.
Furthermore, we conduct additional evaluations of AdaptCRN on the Voicebank+DEMAND dataset, using DeepFilterNet, GTCRN, ULCNet, FSPEN, and LiSenNet for comparison. Results for these models are sourced directly from their respective papers. Given the importance of PESQ as a metric in this evaluation, we ensure a fair comparison by using the results for LiSenNet trained without PESQ loss. As shown in Table VI, the results consistently highlight AdaptCRN’s competitive performance alongside lower computational complexity compared to other models.
@@
VI-C Visualization
A typical audio sample from DNS5 test set is presented in Fig. 3, showcasing noisy and clean spectrograms, enhanced results from DPCRN-light, DPCRN-light integrated with adaptive convolution, and AdaptCRN. This sample poses challenges due to highly non-stationary noise and the presence of two speakers with distinct vocal characteristics, both of which should be preserved in single-channel SE task. The enhanced spectrogram of DPCRN-light exhibits noticeable distortion in one speaker’s voice (highlighted in the green box) and residual noise (highlighted in the blue box), which are effectively mitigated with the incorporation of adaptive convolution. AdaptCRN also achieves superior enhancement quality with ultra-low computational complexity, further highlighting the advantage of the propose methods. Additional audio samples are available at https://github.com/Dahan-Wang/Adaptive-Convolution-for-CNN-based-Speech-Enhancement-Models.
The frame-level kernel attention weights are visualized in Fig. 3(f), demonstrating the relationship between kernel allocation and speech spectral features. The visualization is derived from the third layer of the decoder in DPCRN-light with adaptive convolution, depicting the attention weights of 8 candidate kernels across frames. For clarity, the order of the kernels is manually adjusted during plotting. The results show that the first candidate kernel is predominantly activated during the utterance of the high-pitched speaker, while the segments from the other speaker with a lower pitch are primarily associated with the 2nd to 5th kernels. The 6th to 8th kernels appear to function mainly in pure noise segments. These findings indicate that the adaptive convolution assigns appropriate candidate kernels to each frame based on speech characteristics, effectively facilitating feature extraction and reconstruction.
@@ VII Conclusion
In this paper, we propose adaptive convolution, an efficient and versatile convolutional module for CNN-based SE models. It is a frame-wise causal dynamic convolution that adjusts kernels based on spectral characteristics, akin to adaptive filtering. The adaptive convolution kernel is generated by assembling multiple parallel candidate kernels through frame-level attention weights derived from input features. We introduce several lightweight attention mechanisms to guide the kernel aggregation, which can be extended to a multi-head form for joint multi-layer kernel attention and joint channel attention. Moreover, we present AdaptCRN, an ultra-lightweight model that integrates adaptive convolution and several effective strategies. Experimental results and visualizations demonstrate that adaptive convolution can assign appropriate kernels to each frame based on speech characteristics, thereby significantly improving the performance of multiple CNN-based models with a negligible increase in computational complexity, especially for lightweight models. The results also show that AdaptCRN achieves competitive performance with several other SOTA lightweight models, even those with significantly higher computational cost, which further underscores the efficiency of adaptive convolution and the effectiveness of the proposed lightweight designs.

### v2

@@ VI-A2 Detailed configuration of kernel generation
Adaptive 22449.07 8235.08 15.208 0.777 2.526 3.027 3.309 4.053
TABLE IV: Ablation study results for AdaptCRN on DNS5 test set.
  Para. (K) MACs (M) SI-SNR ESTOI PESQ DNSMOS
OVRL SIG BAK
Noisy - - 5.366 0.625 1.391 2.147 2.918 2.340
AdaptCRN 134.51 40.80 14.892 0.759 2.387 2.939 3.219 4.037
w/o Joint channel attention 124.68 40.18 14.718 0.756 2.352 2.925 3.207 4.031
w/o Joint multi-layer kernel attention 169.21 43.19 14.803 0.759 2.386 2.939 3.221 4.036
w/o Dynamic range compression 134.51 40.80 14.695 0.755 2.342 2.929 3.211 4.031
DW(8)-PW(8)-skip-PW(8) 134.51 40.01 14.672 0.755 2.358 2.931 3.212 4.034
DW(8)-PW(8) 112.88 32.06 14.551 0.754 2.343 2.926 3.207 4.031
DW(8)-PW(64) 255.00 41.41 14.826 0.759 2.390 2.942 3.221 4.039
@@ VI-A3 Generalization capability of adaptive convolution
However, as the model size increases, the improvements introduced by adaptive convolution become less pronounced. For DPCRN-middle and DCCRN, it achieves approximately a 0.06 PESQ gain and a 0.03 DNSMOS-OVRL gain. These improvements are relatively smaller compared to those observed in lightweight models. For DPCRN-large, the improvements are minimal, with only about a 0.2 dB SI-SNR gain, a 0.02 PESQ gain, and a 0.02 DNSMOS-OVRL gain. The diminishing improvement can be attributed to the extensive parameter space inherent in larger models. Specifically, the convolutional kernels in such models already include filters capable of adapting to diverse spectral features, reducing the benefit of additional candidate kernels.
Moreover, as the network size increases, the additional storage and computational overhead introduced by adaptive convolution becomes more substantial. This is primarily due to the increase in the overall size of the convolution kernels in larger models. The impact is particularly pronounced in the large DCCRN model, where convolution operations account for over 97% of the total MACs. In this case, the additional cost introduced by adaptive convolution becomes prohibitively high, rendering the method less suitable for these architectures. These findings indicate that the applicability and effectiveness of the proposed adaptive convolution are limited in large-scale models.
@@ VI-B Experimental Results for AdaptCRN
AdaptCRN 134.51 40.80 14.892 0.759 2.387 2.939 3.219 4.037
Figure 3: Violin plots of (a) SI-SNR, (b) PESQ, and (c) DNSMOS-OVRL results for GTCRN and AdaptCRN on the DNS5 test set.
@@ VI-B1 Ablation study results for AdaptCRN
We evaluate the effectiveness of the proposed structures in AdaptCRN through ablation experiments by selectively removing specific modules. The results, summarized in Table IV, highlight the pivotal role of adaptive convolution in enhancing model performance. With an increase in computational complexity of approximately 7M MACs, adaptive convolution delivers significant improvements, including nearly 1 dB in SI-SNR, 0.02 in ESTOI, 0.19 in PESQ, and over 0.06 in DNSMOS-OVRL. In the proposed model, we utilize adaptive convolution with $K=8$, as this configuration provides a favorable balance between performance gains and parameter amount. Due to the extremely low parameter count of the basic model, the additional parameters introduced by adaptive convolution remain within an acceptable range for real-time applications.
@@
Table IV further presents the model’s performance when the activation function between the two PW Conv layers in the adaptive block is removed (DW(8)-PW(8)-skip-PW(8)), where the numbers in parentheses (e.g., 8) denote $K$, the number of candidate kernels. This modification results in some performance degradation but still outperforms a single PW Conv layer with $K=8$ (DW(8)-PW(8)). This indicate that cascading two PW convolution layers with adaptive convolution is not equivalent to a single layer, unlike in the case of vanilla convolution. However, as analyzed in Section IV-C, this configuration can be reparameterized into a single layer with 64 candidate kernels, showcasing the enhanced nonlinear representational capacity of adaptive convolution. Moreover, directly using 64 independent kernels (DW(8)-PW(64)) achieves better performance compared to the reparameterized result, due to the lower-rank kernel space of the latter. Although this configuration achieves performance comparable to the original AdaptCRN, it requires a larger number of parameters. This result underscores the architectural efficiency of AdaptCRN, which achieves similar performance with fewer redundant parameters.
@@ VI-B2 Comparison with the baseline models
We evaluate our ultra-lightweight model, AdaptCRN, against DPCRN-light, GTCRN, and LiSenNet on DNS5 test set. As illustrated in Table V, AdaptCRN outperforms the baseline models across all metrics. AdaptCRN remains efficient with computational complexity approximately 40M MACs, which is only slightly higher than GTCRN. While it does have more parameters compared to the baselines, it remains within acceptable limits for practical applications. Considering that GTCRN with adaptive convolution has a parameter count and computational complexity similar to AdaptCRN, its results are also presented in Table V for comparison. Despite incorporating adaptive convolution, GTCRN performs significantly worse than AdaptCRN, highlighting the effectiveness of the additional structural designs introduced in AdaptCRN. We further present the SI-SNR, PESQ, and DNSMOS-OVRL results of AdaptCRN and the baseline model with comparable computational complexity, GTCRN, using violin plots in Fig. 3. The results demonstrate that AdaptCRN achieves consistent improvements across all utterances in the test set.
Furthermore, we conduct additional evaluations of AdaptCRN on the Voicebank+DEMAND dataset, using DeepFilterNet, GTCRN, ULCNet, FSPEN, and LiSenNet for comparison. Results for these models are sourced directly from their respective papers. Given the importance of PESQ as a metric in this evaluation, we ensure a fair comparison by using the results for LiSenNet trained without PESQ loss. As shown in Table VI-B, the results consistently highlight AdaptCRN’s competitive performance alongside lower computational complexity compared to other models.
@@
Figure 5: The proportion of frames in which the $k$-th candidate kernel is selected as the dominant kernel across speech or non-speech frames, measured from the third decoder layer of DPCRN-light with adaptive convolution.
VI-C Audio sample visualization
A typical audio sample from DNS5 test set is presented in Fig. VI-B, showcasing noisy and clean spectrograms, enhanced results from DPCRN-light, DPCRN-light integrated with adaptive convolution, and AdaptCRN. This sample poses challenges due to highly non-stationary noise and the presence of two speakers with distinct vocal characteristics, both of which should be preserved in single-channel SE task. The enhanced spectrogram of DPCRN-light exhibits noticeable distortion in one speaker’s voice (highlighted in the green box) and residual noise (highlighted in the blue box), which are effectively mitigated with the incorporation of adaptive convolution. AdaptCRN also achieves superior enhancement quality with ultra-low computational complexity, further highlighting the advantage of the propose methods. Additional audio samples are available at https://github.com/Dahan-Wang/Adaptive-Convolution-for-CNN-based-Speech-Enhancement-Models.
VI-D Correlation between kernel allocation and speech spectral features
We observe that the allocation of candidate kernels in the trained adaptive convolution exhibits a strong correlation with speech spectral features. To illustrate this, the frame-level kernel attention weights are visualized in Fig. VI-B(f). The visualization is derived from the third layer of the decoder in DPCRN-light with adaptive convolution, depicting the attention weights of 8 candidate kernels across frames. For clarity, the order of the kernels is manually adjusted during plotting. The results show that the first candidate kernel is predominantly activated during the utterance of the high-pitched speaker, while the segments from the other speaker with a lower pitch are primarily associated with the 2nd to 5th kernels. The 6th to 8th kernels appear to function mainly in pure noise segments.
We further conduct a quantitative analysis of this relationship on the DNS5 test set. Specifically, we examine the weights assigned to candidate kernels by the kernel attention mechanism of the adaptive convolution layer. For each frame, the candidate kernel with the highest weight is defined as the dominant kernel, as it contributes most strongly to the final convolution kernel. We then compute the proportion of frames in which the $k$-th kernel is selected as the dominant kernel, separately for speech and non-speech frames, with an energy-based voice activity detection (VAD) method used to distinguish between the two categories. The results are shown in Fig. 5. As observed, in speech frames, all candidate kernels have a noticeable chance of being selected as the dominant kernel, with differences across kernels present but not pronounced. In contrast, in non-speech frames, the 6th-8th kernels are significantly more likely to dominate. This suggests that the last three kernels specialize in processing noise-related features, which is consistent with our visualization results. Notably, these kernels also make substantial contributions in speech frames. Two factors may explain this phenomenon: (1) speech frames still contain noise components, necessitating the involvement of noise-related kernels; and (2) VAD often labels short pauses between words as speech frames, meaning that some “speech frames” in practice include non-speech segments, where the last three kernels are more likely to dominate.
These findings indicate that the adaptive convolution assigns appropriate candidate kernels to each frame based on speech characteristics, effectively facilitating feature extraction and reconstruction.
@@ VII Conclusion
In this paper, we propose adaptive convolution, an efficient and versatile convolutional module for CNN-based SE models. It is a frame-wise causal dynamic convolution that adjusts kernels based on spectral characteristics, akin to adaptive filtering. The adaptive convolution kernel is generated by assembling multiple parallel candidate kernels through frame-level attention weights derived from input features. We introduce several lightweight attention mechanisms to guide the kernel aggregation, which can be extended to a multi-head form for joint multi-layer kernel attention and joint channel attention. Moreover, we present AdaptCRN, an ultra-lightweight model that integrates adaptive convolution and several effective strategies. Experimental results and visualizations demonstrate that adaptive convolution can assign appropriate kernels to each frame based on speech characteristics, thereby significantly improving the performance of multiple CNN-based models with a limited increase in computational complexity, especially for lightweight models. The results also show that AdaptCRN achieves competitive performance with several other SOTA lightweight models, even those with significantly higher computational cost, which further underscores the efficiency of adaptive convolution and the effectiveness of the proposed lightweight designs.